# INQUIRE-Search: A Framework for Interactive Discovery in Large-Scale Biodiversity Databases

### v3

@@ Abstract
  1. 1.
    Community science platforms such as iNaturalist contain hundreds of millions of biodiversity images that capture ecological context on behaviors, interactions, phenology, and habitat. Yet ecological workflows to surface these data rely on metadata filtering or manual inspection, leaving this "secondary information" largely inaccessible at scale.
  2. 2.
    We introduce INQUIRE-Search, an open-source system that uses natural language to enable scientists to rapidly search within an ecological image database for specific concepts, verify and export relevant observations, and use these outputs for downstream scientific analysis. Compared to traditional methods, INQUIRE-Search takes a fraction of the time, opening up new possibilities for scientific questions that can be explored.
  3. 3.
    Through five case studies, we demonstrate the range of applications INQUIRE-Search can support, from seasonal variation in behavior across species to forest regrowth after wildfires. These examples demonstrate a new paradigm for interactive, efficient, and scalable scientific discovery that can begin to unlock previously inaccessible scientific value in large-scale biodiversity datasets.
  4. 4.
    Finally, we highlight how AI-enabled discovery tools for science require reframing aspects of the scientific process, including experiment design, data collection, survey effort, and uncertainty analysis.
@@ 1 Introduction
1.1 The Need for Tools to Make Ecological Databases Discoverable
Image data have become a key component of ecological research, collected to document and analyze biodiversity, species interactions, and environmental changes across spatial and temporal scales [oliver2023camera, kays2020born]. With the widespread use of digital photography and automated sensors, community science platforms like iNaturalist [iNaturalist2024] now contain hundreds of millions of image-based species observations [callaghan2021three].
Currently, the scientific use of such databases has primarily focused on species occurrences [kumar2019using, humphreys2019seasonal], but many ecological questions, such as biotic and abiotic interactions, resource use, and habitat associations, require information that extends beyond species identification and location alone. We refer to this contextual information as secondary data: ecologically relevant features that extend beyond the primary observation [pernat2024overcoming, marques2024retrieving, davison2025automated]. Images frequently capture organisms in their microhabitats, with co-occurring species, observable behaviors, or signs of disturbance.
[…]
@@ 1.1 The Need for Tools to Make Ecological Databases Discoverable
1.2 Utilizing Machine Learning to Process Large Scale Data
Recent advances in computer vision provide valuable tools for dataset curation and filtering, and automated species detection models are widely used in ecology [norouzzadeh2018automatically, willi2019identifying]. However, supervised learning approaches to train these models require large, well-labeled datasets, which are often hard to acquire for ecological information of interest. This closed-set design limits the ability of supervised models to generalize beyond fixed categories and constrains their use in open-ended ecological exploration.
Vision-language models (VLMs) offer a powerful alternative. Trained on large image-text datasets, these models map images and language into a shared high-dimensional representation space where semantically similar concepts lie close together [radford2021learning]. For example, an image of a hermit crab wearing plastic as a shell would be embedded into this space closer to the corresponding text "a hermit crab with plastic trash as a shell," and far from unrelated descriptions such as "a yellow car," as seen in Fig 2. VLMs therefore enable us to search for relevant images in a database through a simple similarity search—ranking each image in the database based on its distance to the embedding of a text query.
Figure 2: (Top) Joint vision-language representation space. VLMs are trained to learn a joint embedding space between vision and text modalities, where similar semantic concepts are stored closer together. (Bottom) Image-text search. At search-time, the input text query is embedded by the text encoder of the VLM. Then the query embedding vector is compared against the pre-computed iNaturalist image embeddings to retrieve the most relevant images.
Embedding-enabled search with VLMs has already shown promise for ecological applications. WildCLIP [gabeff2024wildclip] fine-tuned a VLM on camera trap images and ecology-specific language descriptions. From training on large data, this model was able to generalize from concepts seen during training to search large wildlife datasets for new behaviors or habitat details. The INQUIRE Benchmark [vendrow2024inquire] evaluates state-of-the-art VLMs on expert-level open-ended ecological queries across five million iNaturalist 2024 images. Both works clearly demonstrate the potential of such an approach to unlocking new knowledge in databases. Realizing this potential requires a system that connects the broad capabilities of VLMs with the practical needs of ecological research, enabling scientists to make effective use of large image repositories.
To address the need, we introduce INQUIRE-Search, a system for efficient, expert-driven retrieval from large ecological image datasets. As illustrated in Fig 1, the system facilitates an iterative cycle in which researchers pose questions in plain language, retrieve semantically relevant images, assess and refine those results based on ecological validity, and ultimately export curated observations for downstream analysis.
2 Materials and Methods
2.1 INQUIRE-Search System Architecture
INQUIRE-Search is an open-source system for interactive, expert-driven retrieval from large ecological image datasets. It integrates a vision–language embedding model, a high-performance vector index, and a lightweight, browser-based interface to support efficient natural-language search, verification, and data export.
The system uses a state-of-the-art Vision–Language Model (VLM), SigLIP-So400m-384-14 [zhai2023sigmoid], to embed both images and natural-language queries into a shared high-dimensional semantic space. All iNaturalist [iNaturalist2024] images were preprocessed to a uniform format and embedded using the SigLIP image encoder. These embeddings were stored in a FAISS (Facebook AI Similarity Search) [douze2025faiss] index, enabling sub-second similarity search across approximately 300 million images.
At search time, a user-provided text query is embedded and used to rank images in the FAISS index by cosine similarity. The browser-based interface supports metadata filtering (taxonomy, geography, and date), rapid inspection and expert verification of retrieved images, and export of curated observations with complete iNaturalist metadata. Verified image sets can be exported as CSV files, including observation IDs, coordinates, timestamps, taxonomic information, and file URLs, for downstream analysis. Full implementation details are provided in the Appendix.
2.2 INQUIRE-Search Pipeline Overview
We follow a standardized human-in-the-loop workflow using INQUIRE-Search.
  1. 1.
    Query and Prioritize: We composed natural-language queries tailored to each ecological question and applied optional metadata filters (species, date, location). All prompts are listed in Table 1, with complete information on the filters, number of retrievals inspected, and number of samples marked for analysis provided in the Appendix.
  2. 2.
    Retrieve and Verify: For each query, the system returned top-ranked images. Depending on the case study, we screened between 200 and 500 top retrievals for relevant data. A reviewer examined each image and labeled it informative only if it met the case-specific criteria.
  3. 3.
    Export and Analyze: Informative images and their metadata were exported as CSV files for downstream analysis. Analytical procedures were specific to each case study and are described in Section 2.3.
@@ 2.2 INQUIRE-Search Pipeline Overview
Seasonal variation in bird diets Prompts of the form "<species> with <diet type> in its mouth" using: Species (common names): Gray-cheeked Thrush, Ancient Murrelet, American Tree Sparrow, Red-bellied Woodpecker, American Robin Diet types: invertebrate, vertebrate, seed, fruit, nectar and pollen, carrion or animal decay, other plant matter Total: 5 species × 7 diet types = 35 prompts.
Post-fire forest regrowth "Young coniferous trees in burned forest"
"Young deciduous trees in burned forest".
Wildlife mortality "Dead bird"
Plant Phenology "Milkweed germinating or emerging"
"Milkweed flowering"
"Milkweed producing seeds"
"Milkweed dying or withering / in senescence".
Whale re-identification "White underside of humpback whale fluke".
@@
2.3 INQUIRE-Search Case Studies
We collaborated with a diverse set of ecological experts to define a series of scientific questions that would potentially be addressable with information hidden within the iNaturalist database. Together, we used INQUIRE-Search to address each of these questions, designing each experiment to test both the usability and utility of the tool and test the limits of what insights the system could help uncover. Each case study explores a distinct ecological question: (1) seasonal variation in bird diets, (2) post-fire forest regeneration, (3) spatio-temporal patterns of wildlife mortality, (4) plant phenology across seasonal cycles, and (5) individual Re-ID in humpback whales. These examples show how natural language-guided image search and analysis can provide a scalable and flexible approach for extracting ecological insights from large, unstructured image datasets.
2.3.1 CS1: Seasonal Variation in Bird Diets
@@
Overview Diet is a fundamental response and effect trait that shapes vertebrate survival, fitness, trophic position, and ecological interactions, and it varies widely across species and environmental conditions [burin2016omnivory, sibly2012energetics, belmaker2012global]. Although large dietary databases exist [wilman2014eltontraits], seasonal diet information remains sparse: in SAviTraits 1.0 [murphy2023SAviTraits], only about 10% of more than 10,000 bird species are recorded as exhibiting seasonal dietary shifts, a pattern likely driven by data gaps rather than true biological invariance. Extracting such information from community-science imagery is also nontrivial: among over 23,000 relevant iNaturalist observations of American Robin, only 44 are tagged “feeding,” and none specify the type of food consumed. This case study evaluates whether INQUIRE-Search can systematically recover feeding events across seasons from community-sourced photographs and reproduce known dietary patterns. We focus on five species – Gray-cheeked Thrush (Catharus minimus), Ancient Murrelet (Synthliboramphus antiquus), American Tree Sparrow (Spizelloides arborea), Red-bellied Woodpecker (Melanerpes carolinus), and American Robin (Turdus migratorius – which have high dietary certainty scores in SAviTraits [murphy2023SAviTraits], providing a robust reference for comparison.
@@
Search Strategy: For each target species, we construct natural language search queries specifying the bird species and diet type (e.g., "American robin with invertebrate in its mouth"), using the dietary classification system used in SAviTraits. This classification system includes seven dietary categories: (1) invertebrate, (2) vertebrate, (3) seed, (4) fruit, (5) nectar and pollen, (6) carrion or animal decay, and (7) other plant matter. Each diet category was queried separately for summer (June–August) and winter (December–February). Taxonomic filters restricted results to the target species.
Verification: An image was considered informative only when a clearly identifiable food item was visible in the bird’s bill. Top 500 images were screened by a reviewer for every species-diet type-season combination.
Analysis: For each combination, we counted the marked images and calculated the proportion of informative images belonging to each diet category. These proportions were then compared qualitatively with published seasonal diet compositions from SAviTraits.
2.3.2 CS2: Post-fire Forest Regrowth

### v5

# INQUIRE-Search: Interactive Discovery in Large-Scale Biodiversity Databases

@@ Abstract
[…]
@@ 1 Introduction
Observations that capture ecological context beyond species occurrence, such as species interactions and spatiotemporal changes in traits or habitat, are the fundamental underpinning of most ecological research and analyses. Significant resources are invested each year into capturing these observations via, e.g., transects or deployed sensors. Digital imagery is an increasingly common mechanism for ecological observation. Photographs and automated camera systems are primarily used to document species occurrence and are widely applied in studies of species distributions and biodiversity patterns (Oliver et al., 2023; Kays et al., 2020). However, each image also records a broader ecological scene, incidentally capturing information about traits, habitat, seasonality, behavior, and interactions. For example, an image captured to document the occurrence of a bird species may also show the individual feeding, allowing seasonal or inter-species patterns in diet to be inferred; images collected to record species presence in recently disturbed landscapes may incidentally document early stages of forest regeneration; and repeated images of flowering or fruiting plants collected as occurrence records can provide fine-scale phenological information. However, these observations occur opportunistically, when organisms, the environmental conditions, and observers coincide in time and space, making it highly difficult to collect them systematically through targeted field studies.
Community science platforms aggregate opportunistic observations at unprecedented spatial and temporal scales. Platforms such as iNaturalist, Pl@ntNet, and the Macaulay Library now host hundreds of millions of image-based observations worldwide (iNaturalist, 2024; Callaghan et al., 2021; Macaulay Library, 2024). These efforts align with broader open-access biodiversity initiatives, including GBIF and Darwin Core–based infrastructures, which have transformed how species occurrence data are shared and reused globally (GBIF, 2019; Wieczorek et al., 2012). These large-scale databases are widely used in ecological research, but most applications rely on species occurrence information to support analyses of distributions, range limits, and broad-scale patterns of relative abundance (Eichholtzer et al., 2025; Song et al., 2025; Bosso et al., 2024; Moore et al., 2024; Wijewardhana et al., 2022; Deshwal et al., 2021) (Kumar et al., 2019; Humphreys et al., 2019). But these unprecedented scales of data correspond to a higher likelihood of also capturing diverse ecological phenomena from diet to phenology, which is often referred to as “secondary data” in the context of these repositories (Pernat et al., 2024; Marques et al., 2024; Davison et al., 2025). Unfortunately, in the current systems these rare and context-dependent events are extremely difficult to surface in sufficient numbers for analysis. As ecological image repositories such as iNaturalist continue to expand, the challenge is no longer the absence of such data, but the inability to efficiently discover and extract this secondary ecological information at scale.
This inaccessibility stems from how observations are indexed and searched: existing filtering approaches rely heavily on basic metadata (e.g. taxonomic identity, location, and time), which describe the occurrence record but not its ecological context. Richer ecological descriptors (e.g. behavior, interactions, habitat, etc.) are optional, inconsistently applied, or absent altogether; fewer than 25% of the over 300 million iNaturalist records contain any additional annotations (iNaturalist, 2024). For example, a recent study identifying predation events involving rove beetles required manually reviewing approximately 48,000 observations to locate just 159 relevant records (Hu et al., 2025).
Deep learning methods are well-posed to speed up the analysis of large-scale data, particularly with regards to detecting and surfacing observations of specific concepts. In ecology, supervised detection and classification models are already widely used to identify species in camera trap and community science imagery (Norouzzadeh et al., 2018; Willi et al., 2019) (de Castelbajac et al., 2025; Eichholtzer et al., 2025). Training such supervised models to identify secondary ecological information—such as interactions, behaviors, or habitat context—require large, exhaustively labeled datasets that capture their concepts of interest. Additionally, the use of these models would require specialized deep learning expertise and significant data access, storage, and computational infrastructure. This limits the use of supervised deep learning to analyze secondary information at scale.
Recent work has explored methods for efficient filtering and retrieval from large, heterogeneous data collections across modalities, including text (Castro et al., 2024), genomics (Rabelo et al., 2025), and audio (Dong et al., 2015). In the context of image data, vision–language models (VLMs) provide a promising path to reducing these access barriers by enabling flexible, efficient search over visual data with open-ended text, for example, letting an ecologist search all of iNaturalist for “American Robin eating a worm.” Trained on large-scale image-text datasets, VLMs learn to map images and natural language into a shared high-dimensional representation space, where semantically related concepts lie close together (Figure 1) (Radford et al., 2021). This capability enables similarity-based retrieval of images depicting various context without requiring predefined labels. Recent work demonstrates the potential of this approach for ecological discovery: WildCLIP (Gabeff et al., 2024) fine-tuned a VLM on camera trap imagery and ecology-specific language, enabling the retrieval of previously unseen behaviors and habitat attributes from large wildlife datasets, while other related work has evaluated general-purpose vision–language models on expert-defined, open-ended ecological queries across millions of community-science images (Vendrow et al., 2024).
Figure 1: VLMs are trained to learn a joint embedding space between vision and text modalities, where similar semantic concepts are stored closer together. At search-time, the input text query is embedded by the text encoder of the VLM. Then the query embedding vector is compared against the pre-computed iNaturalist image embeddings to retrieve the most relevant images.
Connecting the dots between large-scale image databases and well-designed ecological inference requires us to integrate VLMs enabling flexible text-based query over data with methods for efficient retrieval, verification, and analysis. In this work, we introduce INQUIRE-Search, an open-source, expert-driven system that enables efficient search over large ecological image databases using natural-language queries leveraging VLMs. INQUIRE-Search is designed as a discovery-oriented informatics system, treating ecological image analysis as an interactive information-retrieval task rather than a closed-set predictive modeling problem. Our contributions are threefold:
  • We present a scalable retrieval and verification framework for surfacing complex ecological scenarios—such as interactions, behaviors, and life-history events—that are difficult or infeasible to access using metadata-based or closed-set approaches.
  • We introduce an iterative, human-in-the-loop workflow that supports natural-language querying, expert validation of retrieved images, and export of curated observations for downstream ecological analysis.
  • We demonstrate the ecological utility of this approach through diverse case studies spanning taxa, spatial scales, and ecological phenomena.
2 Methods and System Design
2.1 INQUIRE-Search system architecture
INQUIRE-Search is an open-source system for interactive, expert-driven retrieval from large ecological image datasets. It integrates a state-of-the-art VLM, a high-performance vector index, and a lightweight, browser-based interface to support efficient natural-language search, verification, and data export.
The system uses the SigLIP-So400m-384-14 (Zhai et al., 2023) VLM, to embed both images and natural-language queries into a shared high-dimensional semantic space. All iNaturalist (iNaturalist, 2024) images were preprocessed to a uniform format and embedded using the SigLIP image encoder. These embeddings were stored in a FAISS (Facebook AI Similarity Search) (Douze et al., 2025) index, enabling sub-second similarity search across approximately 300 million images.
At search time, a user-provided text query is embedded and used to rank images in the FAISS index by cosine similarity. The browser-based interface supports metadata filtering (taxonomy, geography, and date), rapid inspection and expert verification of retrieved images, and export of curated observations with complete iNaturalist metadata. Verified image sets can be exported as CSV files, including observation IDs, coordinates, timestamps, taxonomic information, and file URLs, for downstream analysis. Full implementation details are provided in the Appendix A.
2.2 INQUIRE-Search workflow
We follow a standardized human-in-the-loop workflow using INQUIRE-Search (Figure 2).
  1. 1.
    Query and Prioritize: We composed natural-language queries tailored to each ecological question and applied optional metadata filters (species, date, location). All prompts are listed in Table 1, with complete information on the filters, number of retrievals inspected, and number of samples marked for analysis provided in the Appendix B.
  2. 2.
    Retrieve and Verify: Each text query is embedded using the vision–language model and used to retrieve a ranked list of images from the vector index. Images are inspected sequentially in descending similarity order. Across all retrieval tasks, images are labeled as informative only when the target ecological phenomenon is visually identifiable and unambiguous. Inspection budgets are defined ahead of time for each retrieval task (typically 200-500 images) to ensure comparable human effort across evaluations. In some cases, inspection terminated early when extended sequences of non-relevant images indicated declining relevance in the ranked results.
  3. 3.
    Export and Analyze: Informative images and their associated metadata are exported in tabular (CSV) format for downstream analysis. For searches that require spatially constrained data collection, standard GIS operations such as coordinate-based filtering of iNaturalist latitude-longitude metadata and, where necessary, assignment to externally defined spatial boundaries, for filtering and aggregating data. These geographic operations provide spatial context for retrieved datasets and support descriptive analysis rather than formal spatial inference. Task-specific analytical procedures are described in Section 3.
@@ 2.2 INQUIRE-Search workflow
Seasonal variation in bird diets Prompts of the form “<species> with <diet type> in its mouth” using: Species (common names): Gray-cheeked Thrush, Ancient Murrelet, American Tree Sparrow, Red-bellied Woodpecker, American Robin Diet types: invertebrate, vertebrate, seed, fruit, nectar and pollen, carrion or animal decay, other plant matter Total: 5 species × 7 diet types = 35 prompts.
Post-fire forest regrowth “Young coniferous trees in burned forest”
“Young deciduous trees in burned forest”
Wildlife mortality “Dead bird”
Plant Phenology “Milkweed germinating or emerging”
“Milkweed flowering”
“Milkweed producing seeds”
“Milkweed dying or withering / in senescence”
Whale re-identification “White underside of humpback whale fluke”
@@
2.3 Benchmarking and evaluation
For each retrieval task corresponding to a specific scientific question, we evaluate INQUIRE-Search along two complementary directions: 1) retrieval efficiency and 2) ecological utility.
Retrieval efficiency is evaluated using standardized, effort-oriented metrics appropriate for open-ended discovery. Unlike closed-set classification tasks, the ecological discovery problems addressed in this paper lack exhaustive ground-truth labels or known prevalence of relevant events. Under these conditions, traditional accuracy-based metrics such as precision, recall, or mean average precision are not only infeasible but potentially misleading. For each search, we record the number of retrieved images manually inspected ($N_{\text{insp}}$), the number of images verified as informative and retained for analysis ($N_{\text{ret}}$; final dataset size), and the screening yield ($Y=N_{\text{ret}}/N_{\text{insp}}$). It is important to note that screening yield is not interpreted as a measure of model performance or retrieval accuracy, since the prevalence of relevant observations varies across queries and study contexts, but is used comparatively to assess how efficiently INQUIRE-Search concentrates relevant observations under comparable human effort for inspection.
To enable such comparison, we evaluate INQUIRE-Search against a traditional iNaturalist filtering baseline representing best-practice metadata-based search. Baseline queries apply the same taxonomic, spatial, and temporal filters used in INQUIRE-Search to ensure comparable search scope. Beyond these shared constraints, baselines rely on structured metadata fields when relevant controlled attributes exist, and otherwise use keyword matching over unstructured text fields such as observation descriptions, tags, and annotations. Keyword queries are constructed to approximate the semantic intent of the corresponding INQUIRE-Search prompts as closely as possible. All retrieved results are screened using identical expert verification criteria. Where applicable, we report the yield ratio ($Y_{\text{ratio}}=Y_{\text{INQUIRE-Search}}/Y_{\text{baseline}}$), which captures the relative efficiency between INQUIRE-Search and the baseline method under fixed inspection budgets. All keyword queries are specified within the Results of each case study.
Ecological utility is assessed by examining whether the datasets produced by INQUIRE-Search for a given retrieval task are sufficient to support meaningful downstream ecological analyses. These analyses are described in Section 3.
2.4 Data availability and reproducibility
INQUIRE-Search is released as an open-source system, and the full codebase, documentation, and instructions for reproducing all experiments are available at https://github.com/Beery-Lab/INQUIRE-Search. Case study datasets are derived from openly accessible iNaturalist observations and are exported with complete associated metadata (including observation identifiers, coordinates, timestamps, taxonomic information, and licenses) to support transparency and reproducibility.
3 Validation with Ecological Case Studies
To illustrate how this methodology could support ecological analyses, we present a set of preliminary, scientifically diverse case studies. We first identified questions that require contextual information—such as behaviors, interactions, or life-history events. We focus on phenomena that are likely to be present and visible in community-science images but are rarely annotated, and therefore difficult to access at scale. Each case study is designed to evaluate the ability of INQUIRE-Search to retrieve relevant data, rather than to conduct a full ecological study. Using INQUIRE-Search, we recovered and analyzed relevant observations for each question, designing each experiment to evaluate the usability, flexibility, and limitations of the tool rather than to produce definitive scientific conclusions. Each case study explores a distinct question: (1) seasonal variation in bird diets, (2) post-fire forest regeneration, (3) spatio-temporal patterns of wildlife mortality, (4) plant phenology across seasonal cycles, and (5) individual Re-ID in humpback whales. These examples demonstrate the validity and diversity of carefully-designed natural language-guided image search as a novel mechanism for data collection, providing a scalable and flexible approach for extracting ecological insights from large, unstructured image datasets.
3.1 CS1: Seasonal variation in bird diets
Overview. This case study evaluates whether INQUIRE-Search can recover feeding events across seasons from community-sourced photographs and reproduce known dietary patterns. We compare the trends in INQUIRE-Search retrieved dataset against published dietary records in SAviTraits (Murphy et al., 2023).
@@ 3.1 CS1: Seasonal variation in bird diets
Diet is a fundamental response and effect trait that shapes vertebrate survival, fitness, trophic position, and ecological interactions, and it varies widely across species and environmental conditions (Burin et al., 2016; Sibly et al., 2012; Belmaker et al., 2012). Although large dietary databases exist (Wilman et al., 2014), seasonal diet information remains sparse: in SAviTraits 1.0 (Murphy et al., 2023), only about 10% of more than 10,000 bird species are recorded as exhibiting seasonal dietary shifts. This reflects the difficulty of documenting diet across seasons: seasonal feeding requires repeated observations across time, often during unpredictable windows, and is typically inferred from labor-intensive field studies or scattered literature reports. Community-science image repositories may fill these gaps by capturing feeding events as secondary data, including seasons and contexts that are difficult to sample systematically through targeted field studies. We try to surface these observations efficiently using INQUIRE-Search. In this case study, we focus on five species—Gray-cheeked Thrush (Catharus minimus), Ancient Murrelet (Synthliboramphus antiquus), American Tree Sparrow (Spizelloides arborea), Red-bellied Woodpecker (Melanerpes carolinus), and American Robin (Turdus migratorius)—which have high dietary certainty scores in SAviTraits (Murphy et al., 2023), providing a robust reference for comparison.
@@
INQUIRE-Search workflow for CS1. 1) Query and Prioritize: For each target species, we construct natural language search queries specifying the bird species and diet type (e.g., “American robin with invertebrate in its mouth”), using the dietary classification system used in SAviTraits. This classification system includes seven dietary categories: (1) invertebrate, (2) vertebrate, (3) seed, (4) fruit, (5) nectar and pollen, (6) carrion or animal decay, and (7) other plant matter. Each diet category was queried separately for summer (June–August) and winter (December–February). Taxonomic filters restricted results to the target species. 2) Retrieve and Verify: The top 500 images for each species-diet type-season combination were screened by a reviewer to filter for images with a clearly identifiable food item visible in the bird’s bill, with maximum of $N_{insp}=7000$ per species (some species had less then 500 observations for a specific season, resulting in smaller $N_{insp}$). 3) Export and Analyze: For each combination, we counted the marked images and calculated the proportion of informative images belonging to each diet category to compare qualitatively with published seasonal diet compositions from SAviTraits. When comparing retrieval across species, we compare $N_{ret}$.
Results. Across the five study species, INQUIRE-Search revealed substantial variation in the availability of dietary information in community-science platforms (Table 2). Aggregated across diet categories and seasons, INQUIRE-Search retained the highest number of informative images ($N_{ret}=669$) for American Robin, followed by Red-bellied Woodpecker and American Tree Sparrow ($N_{ret}=395$ and $188$, respectively), while for both Ancient Murrelet and Gray-cheeked Thrush we found a negligible number of images containing dietary information ($N_{ret}<=2$).
Notably, standard iNaturalist search tools recovered markedly fewer diet observations. Keyword-based queries (i.e. species filtering and using “eating” + diet term as a keyword search) returned $N_{ret}=105$ feeding observations for American Robin, $27$ for Red-bellied Woodpecker, and $12$ for American Tree Sparrow, with zero results for the remaining species. In many cases, baseline queries returned fewer total images than the inspection budget, indicating that diet-related feeding events are sparsely indexed in structured metadata and text fields. In practice, these are insufficient for seasonal comparison.
Figure 3: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for “Red-bellied Woodpecker” and “Invertebrate.” (Bottom) Diet comparisons. INQUIRE-Search yields dietary patterns that align closely with those documented in the SAviTraits reference dataset.
Where sufficient data were available, INQUIRE-Search results closely match known seasonal diet compositions reported in SAviTraits, as seen in Figure 3, with agreement in both dominant food types and their relative compositions. For example, American Robin summer diets were composed by roughly equal proportions of fruits and invertebrates, while Red-bellied Woodpecker showed a diet dominated by seeds (50% in INQUIRE-Search vs. 40% in SAviTraits). In winter, all three analyzable species exhibited diets strongly dominated by a single food category, reflecting increased specialization during colder months.
These results show that INQUIRE-Search can recover large numbers of verifiable feeding events and reproduce known seasonal diet trends, provided that the focal species is frequently photographed within community science repositories. Two of the species, however, lacked sufficient image data for meaningful analysis despite being well-studied in the scientific literature and represented in databases like SAviTraits. This limitation is likely even more pronounced for understudied species—precisely those for which new dietary data are most needed—because they are also underrepresented in iNaturalist.
3.2 CS2: Post-fire forest regrowth
Overview. This case study evaluates whether INQUIRE-Search can identify young coniferous and deciduous individuals within the 2012 High Park Fire perimeter from community-sourced photographs and examine how their occurrence varies across burn severity classes (Figure 4).